3/5 of a road is to be resurfacedif road 13/5 km long what length in kilometre
 

Grade:6

4 Answers

Ritika Das
6 years ago
The length in kilometre would be 39/25 km if 3/5 of a road is to be resurfaced as 3/5 * 13/5 is the answer. 
Nitesh
6 years ago
Total length of the road is 13/5 kilometre. Resurfaced part of road is 3/5 of total length.So total length of Resurfaced road is 13/5×3/5 Km=39/25 Km
Gautam kumar
5 years ago
3/5 of a road is to be resurfaced.if the road is 13/5 km long.what length in kilometers is to be resurfaced?
Yuvraj Singh
5 years ago
The resurfaced length in km = 39/25.
The fraction of road resurfaced is 3/5 of total of 13/5 km lenth.
So resurfaced length =  3/5*13/5 = 39/25.